A VectorA is added to B=6i-8j the resultant vector is in the positive x direction and has a magnitude equal to A CALCULATE the direction of A
1
Expert's answer
2019-10-14T09:39:45-0400

Let


"\bold{A}=x\bold{i}+y\bold{j}"

"\bold{A+B}=(x+6)\bold{i}+(y-8)\bold{j}=z\bold{i}"

"y=8, x+6=z"

But


"|\bold{A}|=|\bold{A+B}|\to z^2=x^2+y^2"

"(x+6)^2=x^2+8^2"

"x=\frac{7}{3}"

The direction of A is


"\theta=\arctan{\frac{y}{x}}=\arctan{\frac{8}{\frac{7}{3}}}=73.7\degree"

counterclockwise from the positive x direction.